OpenStudy (bibby):

Determine if the series converges or diverges. \[\sum_{n=2}^{∞} \frac{4n^{2}-5}{3n^3+2n}\]

OpenStudy (bibby):

\[\sum_{n=2}^{∞} \frac{4n^{2}-5}{3n^3+2n}\]

OpenStudy (anonymous):

Compare to the divergent series, \(\displaystyle \sum_{n=2}^\infty \frac{1}{n}\)

OpenStudy (bibby):

How do you decide that? is it because the highest order exponents simplify to 4/3n?

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

is this a scam

OpenStudy (anonymous):

? You just have to show that \[\frac{4n^2-5}{3n^3+2n}>\frac{1}{n}\]

OpenStudy (bibby):

and if you divide out the ns you get 4/3n. How do I prove that \[\frac{4}{3n} > \frac{1}{n}\]

OpenStudy (anonymous):

(4/3) times something is always bigger than that something.

OpenStudy (bibby):

oh wait I am severely mentally retarded forgive me

OpenStudy (bibby):

thank

OpenStudy (anonymous):

Piss

OpenStudy (anonymous):

yw